value of 2-3*5+2/4+5 is


1
Expert's answer
2021-05-07T10:17:15-0400

 2-3*5+2/4+5="2 - 15 + \frac{1}{2} + 5 = 2 - 15 + 5 + \frac{1}{2} = - 8 + 0.5 = - 7.5"

Answer:-7.5